OpenStudy (anonymous):

Solve the equation, Express radicals in simplest form, 4over9w^2-4over3w=-1

OpenStudy (anonymous):

is it ((2/(3w) -1)^2 =0 is w=2/3 or is it ((2w/(3) -1)^2 =0 is w=3/2